OpenStudy (anonymous):

QUIZ TO MEDAL AND FAN!!! IN 1 MINUTE

OpenStudy (anonymous):

whats the Q?

OpenStudy (anonymous):

|dw:1400280615007:dw qUESTION 1, with a rectangle with sides; L= 20 W=34 H= 18 what is the surface area?|

OpenStudy (anonymous):

that's it actuLLY

OpenStudy (anonymous):

3304

OpenStudy (anonymous):

i dont see the picture

OpenStudy (anonymous):

:1400280615007:dw

OpenStudy (anonymous):

1400280615007:dw

OpenStudy (anonymous):

nope no matter how its posted i cant see it

OpenStudy (anonymous):

U DONT NEED THE PIC srry caps

OpenStudy (anonymous):

@eragondragon666 its simple surface area formual: A=2(WL+HL+HW) plug in numbers: A=2(34*20+18*20+18*34) do math and get answer: A=3304

OpenStudy (anonymous):

okay but dont give the answers till you explain as in post that description not just a=3304 k?

OpenStudy (anonymous):

i know that lol it was just very easy so yea lol